OpenStudy (anonymous):

Which equation corresponds to the sentence, "Five times a minus two is equal to t"?

OpenStudy (anonymous):

t=-2*5

OpenStudy (anonymous):

Hmm..somehow, I interpret that as 5*(a-2) = t

OpenStudy (anonymous):

Ah, I misread. I believe you're correct.

jhonyy9 (jhonyy9):

5a-2=t